OpenStudy (anonymous):

Determine the acceleration of a car starting from rest and increasing in speed to 60 m/s in 15 seconds

jhonyy9 (jhonyy9):

900 m/s squared

OpenStudy (anonymous):

Can you show your steps please?

jhonyy9 (jhonyy9):

do you know the formula of acceleration ?

jhonyy9 (jhonyy9):

a=v*t

OpenStudy (anonymous):

Yeah. I kept using s = d/t and it didn't work

jhonyy9 (jhonyy9):

yes because there in this exersice hav not added those terms

OpenStudy (anonymous):

I see now . Thanks

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (cornzyblack):

Or you can use this formula v= u + at v= final velocity which is 60m/s u= initial velocity which is 0m/s t= time which is 15seconds a= acceleration if we make a the subject of formula then, a= (v - u)/ t so a= 4m/s^2